Graph the line with the slope of -4/3 that passes through the points(5,4) Then find the equation
Answer by mananth(16946)

Graph the line with the slope of -4/3 that passes through the points(5,4) Then find the equation
...
m=-4/3
point (5,4)
first find b in the equation
y=mx+b
plug x,y,m
4=-4/3*5 +b
4=-20/3+b
4+(20/3)= b
12+20 / 3=b
32/3 =b
...
equation
y= -4x/3 +32/3
...
m.ananth@hotmail.ca